# A unified approach to hypergeometric class functions

### v1

@@ 1 Introduction
In the literature one can find several names for this family. In this paper, we will use the name hypergeometric class equations for equations of the form (1.1) satisfying (1.2).111 In the book by Nikiforov-Uvarov [NU], and also in [De1, De2], this family is called equations of the hypergeometric type. However, Slavianov-Lay’s book [SL] suggests to use the term “type” for smaller families, such as the families (1)–(5) listed below. [DIL] calls this family grounded hypergeometric class. The name hypergeometric class, following the suggestions of Slavianov-Lay, is reserved in [DIL] for a somewhat wider family, where $\eta(z)$ is allowed to be a rational function such that $\eta\sigma$ is a polynomial of degree $\leq 2$. Anyway, in this paper for simplicity we drop the adjective “grounded”.
@@
The unified approach has its limitations. There are some properties that are not easy to formulate in a unified way. For instance, only one pair of recurrence relations depends analytically on the coefficients. If we restrict ourselves to specific types, then we often find a bigger number of recurrence relations (eg. at least 12 in the case of the Gauss hypergeometric equation). Another family of important properties not included in our presentation are quadratic identities, which link various types of hypergeometric class functions. Thus this work should be compared with other studies of hypergeometric class equations, such as [De1], where specific types are described one by one.
@@
which will be called the Casimir operator of Miller’s Lie algebra. In the representation (1.4), restricted to the quadric 1.5 and the eigenspace $N=0$, we obtain a 2nd order differential operator in $z$, denoted ${\mathcal{C}}(\sigma,\kappa)$. Its eigenvalue equation
$\big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\big{)}f=0$ (1.7)
@@
We devote Section 3 to the properties of hypergeometric class equations and functions which follow directly from the properties of Miller’s Lie algebra: the basic recurrence relations and the basic symmetry. Thanks to the recurrence relations, once we know a solution of a certain hypergeometric class equation, we know its solutions for a whole ladder of parameters labelled by $n\in{\mathbb{Z}}$.
@@
Certain ladders are special. One of them contains solutions which can be expressed in terms of elementary functions. We call them Chebyshev solutions, because it contains the well-known Chebyshev polynomials. It is described in Subsection 3.6.
@@
Another ladder of solutions consists of polynomials. In the convention that we adopted, this ladder is descending—its elements are zero for $n>0$. We devote to this ladder the whole Section 7. This family includes the famous classical orthogonal polynomials. The basic symmetry produces from the polynomial ladder an ascending ladder, which needs no separate discussion.
@@
For the Hermite equation the polynomial $\sigma$ does not have a zero. Other types, that is equations reducible to the ${}_{2}F_{1}$, ${}_{1}F_{1}$, ${}_{2}F_{0}$ and ${}_{0}F_{1}$ types, $\sigma$ has at least one zero. We devote Section 4 to these equations, where without loss of generality we assum $\sigma(0)=0$. We can then write a formal power series $F(\sigma,\kappa,\omega;z)$ which solves the equation
$\big{(}{\mathcal{C}}(\sigma,\kappa)+\omega\big{)}F=0,$ (1.9)
@@
normalized to be $1$ at zero. This power series is convergent if $\sigma^{\prime}(0)\neq 0$, that is for ${}_{2}F_{1}$, ${}_{1}F_{1}$, and ${}_{0}F_{1}$. If $\sigma^{\prime}(0)=0$, that is for ${}_{2}F_{0}$, it is divergent, but it is asymptotic to a well-defined solution. Thus we obtain a unification of 4 types of hypergeometric functions in a single function $F(\sigma,\kappa,\omega;z)$, which we call the unified hypergeometric function. It depends meromorphically on $5$ complex parameters (two parameters for $\sigma$, two for $\kappa$ and one for $\omega$).
@@
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, it is natural to assume $\sigma^{\prime}(0)=1$ and to normalize the unified hypergeometric function by dividing it by $\Gamma(m+1)$, where $m=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. This is sometimes called Olver’s normalization. Thus we obtain the Olver’s normalized unified hypergeometric function $\mathbf{F}(\sigma,\kappa,\omega;z)$, which depends analytically on $4$ complex parameters (one parameter for $\sigma$, two for $\kappa$ and one for $\omega$).
@@
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$ the hypergeometric class equation has an additional discrete symmetry, which involves the gauge transformation $x^{m}\cdot x^{-m}$. We call it the power symmetry.
@@
If $\sigma(0)=0$, and $\sigma^{\prime\prime}\neq 0$ or $\kappa^{\prime}\neq 0$, that is in the cases of the ${}_{2}F_{1}$, ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations, we have yet another discrete symmetry, which we call the inversion symmetry. For the ${}_{2}F_{1}$ equation it involves the change of the variable $z\mapsto\frac{1}{z}$ and the transformed equation is still of the form ${}_{2}F_{1}$. In the confluent cases it interchanges the ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations, and it involves $z\to-\frac{1}{z}$. It is however not defined for the ${}_{0}F_{1}$ equation.
@@
As we mentioned above, some ladders of solutions of hypergeometric class equations consist of polynomials. These polynomials include the famous classical orthogonal polynomials (sometimes called the very classical orthogonal polynomials). They allow for a very elegant unified treatment, which includes not only recurrence relations, but also generating functions (the famous Rodrigues formula) and the orthogonality relations. All of this is briefly described in Section 7.
As we mentioned above, if $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, that is in the cases equivalent to types ${}_{2}F_{1}$, ${}_{1}F_{1}$, and ${}_{0}F_{1}$, we have the power symmetry. Because of that, these equations have two linearly independent solutions with a distinct behavior at zero: the unified hypergeometric function, which is analytic at zero, and another solution behaving as $\sim z^{m}$, where $m=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. The linear independence breaks down for $m\in{\mathbb{Z}}$, when both solutions are proportional to one another. This case is called degenerate and is discussed in Section 8. The unified hypergeometric function in this case has an additional integral representation and an elegant generating function.
@@
Almost all our discussion is algebraic, without functional analysis. We are aware that it is natural to view hypergeometric class operators as closed (or self-adjoint) Sturm-Liouville operators on appropriate weighted $L^{2}$ spaces. It would be also interesting to consider representations of Miller’s Lie algebra in Hilbert spaces. This would, however, require breaking our discussion into separate types. We use some elements of functional analysis only in Section 7. about classical orthogonal polynomials. We show how to view them as eigenfunctions of certain self-adjoint Sturm-Liouville operators on weighted $L^{2}$ spaces with appropriate boundary conditions—this can be done in a rather unified fashion.
@@
Under the name “equations of the hypergeometric type” they were considerd in a unified way in the book by Nikiforov–Uvarov [NU]. This book was one of the two main inspirations for our article. A part of the material of our work is adapted from [NU], notably the material of Subsections 3.5, 5.1, 7.2, 7.6.
@@
Let us also mention earlier works that contained elements of a unified theory of hypergeometric class equations—[IH] and [Tr].
@@
The family of hypergeometric class polynomials that form an orthogonal basis in an appropriate weighted Hilbert space consists essentially of Jacobi, Laguerre and Hermite polynomials, often called classical orthogonal polynomials It has an especially large literature, e.g. [NU, R]. In the more recent literature the name “classical orthogonal polynomials” is sometimes given to a broader family, given by the so-called Askey scheme. Some authors proposed to use the name very classical orthogonal polynomials for the family consisting of Jacobi, Laguerre and Hermite polynomials.
@@ 2.1 Three low-dimensional Lie algebras
Define the linear map $\tau:m_{\alpha,\beta}\to m_{\alpha,-\beta}$ by
@@
Then $\tau$ is an isomorphism. The identity automorphism together with $\tau$ form a group isomorphic to ${\mathbb{Z}}_{2}$, which will be called the Weyl group of $m_{\alpha,\beta}$.
@@ 2.3 Casimir
Extend the isomorphisms $\tau$ and $\pi_{n}$ defined in (2.21) and (2.22) to the algebra of operators on ${\cal Z}$. Then
$\tau({\mathcal{C}}_{\alpha,\beta})$ $={\mathcal{C}}_{\alpha,-\beta},$ (2.27)
@@ 2.4 Ladders
It is easy to see that if ${\cal Z}$ is irreducible and $N$ possesses an eigenvalue, then the spectrum of $N$ is a ladder. In fact this follows from
@@
For $n\in{\mathbb{C}}$ we define
${\cal Z}^{n}:=\{v\in{\cal Z}\ :\quad\quad Nv=nv\}.$ (2.32)
$A_{\pm}^{n\pm\frac{1}{2}}{\mathcal{C}}^{n}={\mathcal{C}}^{n}A_{\pm}^{n\pm\frac{1}{2}}.$ (2.33)
@@ 2.5 Representation by 1st order differential operators
[…]
@@ 2.6 Implementation of isomorphisms
Let us go back to ${\mathbb{C}}^{3}$. The implementation of the isomorphism $\pi_{n}$ defined in (2.22) is simple:
$\pi_{n}(\cdot)=t^{-n}\cdot t^{n}.$ (2.42)
Let us now describe an implementation of the isomorphism $\tau$. Let $\rho(z)$ solve
@@
$\begin{array}[]{l}T\check{g}(w,z)=\rho(z)\check{g}(\frac{\sigma(z)}{w},z)\end{array}.$
@@
We easily see that in the coordinates (2.39)
$T^{n}F(z):=\rho(z)\sigma^{n}(z)F(z).$
@@ 3 Basic properties of hypergeometric class equations
In this section we introduce hypergeometric class equations. They will be presented as the eigenvalue equation of a certain second order differential operator ${\mathcal{C}}(\sigma,\kappa)$. We discuss a number of properties of these equations that can be described in a unified way: the basic symmetry, basic factorization and basic recurrence relations.
The operator ${\mathcal{C}}(\sigma,\kappa)$ is essentially the Casimir of Miller’s Lie algebra at $N=0$, and all its properties discussed in this section follow from the properties of Miller’s Lie algebra analyzed in the previous section. However, this section can be read independently.
@@ 3.1 Remark about notation
Let $a,b,c$ be certain complex functions. Instead of saying that we consider the equation
@@ 3.2 Parametrization of hypergeometric class operators
where $\sigma,\tau,\eta$ satisfy the conditions (1.2). However, the parametrization of these equations with $\sigma,\tau,\eta$ is not always convenient. In this subsection we introduce another parametrization, which is usually better.
@@
${\cal C}(\sigma,\kappa)$ $:=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa(z)\big{)}\partial_{z}+\tfrac{1}{2}\kappa^{\prime}$ (3.5)
@@
${\cal C}(\sigma,\kappa)+\omega$ (3.6)
Note that ${\mathcal{C}}(\sigma,\kappa)$ coincides with ${\mathcal{C}}^{0}$, the Casimir operator for Miller’s Lie algebra restricted to the subspace $N=0$, see (2.41).
@@ 3.3 Basic symmetry
${\cal C}(\sigma,\kappa)=\rho^{-1}(z)\partial_{z}\sigma(z)\rho(z)\partial_{z}+\tfrac{1}{2}\kappa^{\prime}.$ (3.8)
@@
$\begin{array}[]{l}\rho(z){\cal C}(\sigma,\kappa)\rho^{-1}(z)={\cal C}(\sigma,-\kappa).\end{array}$ (3.9)
@@
${\cal C}(\sigma,\kappa)F=0\ \Rightarrow\ {\cal C}(\sigma,-\kappa)\rho F=0.$ (3.10)
Proof.  We multiply the equation
$0=\left(\rho^{-1}(z)\partial_{z}\sigma(z)\rho(z)\partial_{z}+\tfrac{1}{2}\kappa^{\prime}\right)f(z)$
with $\rho(z)$. We get
$\begin{array}[]{rl}0&=\left(\partial_{z}\sigma(z)\rho(z)\partial_{z}\rho^{-1}(z)+\frac{1}{2}\kappa^{\prime}\right)\rho(z)f(z)\\[8.53581pt] &=(\sigma(z)\partial_{z}^{2}+(-\kappa(z)+\sigma^{\prime}(z))\partial_{z}-\frac{1}{2}\kappa^{\prime})\rho(z)f(z).\end{array}$
@@
Applying twice the basic symmetry we get the identity. Hence we obtain a group of symmetries of hypergeometric class equations isomorphic to ${\mathbb{Z}}_{2}$.
Note that the basic symmetry of Theorem 3.10 corresponds to the symmetry $\tau$ of Miller’s Lie algebra, see (2.21) and Theorem 2.1.

### v2

@@ 1 Introduction
In the literature one can find several names for this family. In this paper, we will use the name hypergeometric class equations for equations of the form (1.1) satisfying (1.2). [^1] In the book by Nikiforov-Uvarov [NU], and also in [De1, De2], this family is called equations of the hypergeometric type. However, Slavianov-Lay’s book [SL] suggests to use the term “type” for smaller families, such as the families (1)–(5) listed below. [DIL] calls this family the grounded Riemann class. The name Riemann class, following the suggestions of Slavianov-Lay, is reserved in [DIL] for a somewhat wider family, where $\eta(z)$ is allowed to be a rational function such that $\eta\sigma$ is a polynomial of degree $\leq 2$. See Appendix B for more comments.
@@
The unified approach has its limitations. There are some properties that are not easy to formulate in a unified way. For instance, only one pair of recurrence relations depends analytically on the coefficients. If we restrict ourselves to specific types, then we often find a bigger number of recurrence relations (eg. at least 12 in the case of the Gauss hypergeometric equation). Another family of important properties not included in our presentation are quadratic identities, which link various types of hypergeometric class functions. Thus this work should be compared to other studies of hypergeometric class equations, such as [De1], where specific types are described one by one.
@@
which will be called the Casimir operator of Miller’s Lie algebra. In the representation (1.4), the operator (1.6) restricted to the quadric 1.5 and to the eigenspace $N=0$ is a 2nd order differential operator in $z$, which we denote ${\cal H}(\sigma,\kappa)$. Its eigenvalue equation
$\big{(}{\cal H}(\sigma,\kappa)+\omega\big{)}f=0$ (1.7)
@@
We devote Section 3 to the properties of hypergeometric class equations and functions which follow directly from the action of root operators and Weyl symmetry of Miller’s Lie algebra: the basic recurrence relations and the basic symmetry. Thanks to the recurrence relations, once we know a solution of a certain hypergeometric class equation, we know its solutions for a whole ladder of parameters labelled by $n\in{\mathbb{Z}}$.
@@
Certain ladders are special. One of them contains solutions which can be expressed in terms of elementary functions. We call it the Chebyshev ladder, because it contains the well-known Chebyshev polynomials. It is described in Subsection 3.6. See also [DGR] Subsec. 3.3 and 4.5, [Poole], Sec. 31.
@@
Another ladder of solutions consists of polynomials. In the convention that we adopted, this ladder is descending—its elements are zero for $n>0$. We devote to this ladder the whole Section 7. The basic symmetry produces from the polynomial ladder an ascending ladder, which needs no separate discussion.
@@
For the Hermite equation the polynomial $\sigma$ does not have a zero. For other types, that is for equations reducible to the ${}_{2}F_{1}$, ${}_{1}F_{1}$, ${}_{2}F_{0}$ and ${}_{0}F_{1}$ types, $\sigma$ has at least one zero. We devote Section 4 to these equations, where without loss of generality we assume $\sigma(0)=0$. We can then write a formal power series $F(\sigma,\kappa,\omega;z)$ which solves the equation
$\big{(}{\cal H}(\sigma,\kappa)+\omega\big{)}F=0,$ (1.9)
@@
normalized to be $1$ at zero. This power series is convergent if $\sigma^{\prime}(0)\neq 0$, that is for ${}_{2}F_{1}$, ${}_{1}F_{1}$, and ${}_{0}F_{1}$. If $\sigma^{\prime}(0)=0$, that is for ${}_{2}F_{0}$, it either terminates and is a polynomial, or is divergent. Nevertheless, even in the divergent case this series is asymptotic to a well-defined function ${}_{2}F_{0}$, which we separately discuss in Appendix A. Thus we obtain a unification of 4 types of hypergeometric functions in a single function $F(\sigma,\kappa,\omega;z)$, which we call the unified hypergeometric function. It depends meromorphically on $5$ complex parameters (two parameters for $\sigma$, two for $\kappa$ and one for $\omega$).
@@
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, it is natural to assume $\sigma^{\prime}(0)=1$ and to normalize the unified hypergeometric function by dividing it by $\Gamma(m+1)$, where $m=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. This is sometimes called Olver’s normalization. Thus we obtain the Olver normalized unified hypergeometric function $\mathbf{F}(\sigma,\kappa,\omega;z)$, which depends analytically on $4$ complex parameters (one parameter for $\sigma$, two for $\kappa$ and one for $\omega$).
@@
If $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, the hypergeometric class has an additional discrete symmetry, which we call the power symmetry. This transformation involves gauging with the function $x^{m}$, that is, $(\cdot)\mapsto x^{m}(\cdot)x^{-m}$. (In the literature this transformation is sometimes called an F-homotopy.)
@@
If $\sigma(0)=0$, and $\sigma^{\prime\prime}\neq 0$ or $\kappa^{\prime}\neq 0$, that is for the ${}_{2}F_{1}$, ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations, we have yet another discrete symmetry, which we call the inversion symmetry. For the ${}_{2}F_{1}$ equation it involves the change of the variable $z\mapsto\frac{1}{z}$ and the transformed equation is still of the form ${}_{2}F_{1}$. In the confluent cases it interchanges the ${}_{1}F_{1}$ and ${}_{2}F_{0}$ equations, and it involves $z\to-\frac{1}{z}$. It is however not defined for the ${}_{0}F_{1}$ equation.
@@
As we mentioned above, some ladders of solutions of hypergeometric class equations consist of polynomials. These polynomials include the famous classical orthogonal polynomials (sometimes called the very classical orthogonal polynomials). They allow for a very elegant unified treatment, which includes not only recurrence relations, but also generating functions, the famous Rodrigues formula and the orthogonality relations. All of this is briefly described in Section 7.
As we mentioned above, if $\sigma(0)=0$ and $\sigma^{\prime}(0)\neq 0$, that is in the cases equivalent to types ${}_{2}F_{1}$, ${}_{1}F_{1}$, and ${}_{0}F_{1}$, we have the power symmetry. Because of that, these equations have two linearly independent solutions with a distinct behavior at zero: the unified hypergeometric function, which is analytic at zero, and another solution behaving as $\sim z^{m}$, where $m=\frac{\kappa(0)}{\sigma^{\prime}(0)}$. The linear independence breaks down for $m\in{\mathbb{Z}}$, when both solutions are proportional to one another. This case is called degenerate and is discussed in Section 8. The Olver normalizzed unified hypergeometric function in this case has an additional integral representation and an elegant generating function.
@@
Almost all our discussion is algebraic, without functional analysis. We are aware that it is natural to view hypergeometric class operators as closed (or self-adjoint) Sturm-Liouville operators on appropriate weighted $L^{2}$ spaces. It would be also interesting to consider representations of Miller’s Lie algebra in Hilbert spaces. This would, however, require breaking our discussion into separate types. The only place where we use some elements of functional analysis is Section 7 about classical orthogonal polynomials. We show how to view them as eigenfunctions of certain self-adjoint Sturm-Liouville operators on weighted $L^{2}$ spaces with appropriate boundary conditions—this can be done in a rather unified fashion.
@@
Under the name “equations of the hypergeometric type” they were considered in a unified way in the book by Nikiforov–Uvarov [NU]. This book was one of the two main inspirations for our article. A part of the material of our work is adapted from [NU], notably the material of Subsections 3.5, 5.1, 7.2, 7.6.
@@
There are many works that contain elements of a unified theory of hypergeometric class equations, the idea of recurrence relations and factorizations. Among early ones let us mention the classic work of Infeld and Hull [IH] and of Truesdell [Tr]. Later treatments include [CKS, SHD].
@@
The family of hypergeometric class polynomials that form an orthogonal basis in an appropriate weighted Hilbert space consists essentially of Jacobi, Laguerre and Hermite polynomials, often called classical orthogonal polynomials. It has an especially large literature, e.g. [NU, R]. In the more recent literature the name “classical orthogonal polynomials” is sometimes given to a broader family, given by the so-called Askey scheme. Some authors proposed to use the name very classical orthogonal polynomials for the family consisting of Jacobi, Laguerre and Hermite polynomials.
@@ 2.1 Three low-dimensional Lie algebras
Define the linear map ${\varepsilon}:m_{\alpha,\beta}\to m_{\alpha,-\beta}$ by
@@
Then ${\varepsilon}$ is an isomorphism. The identity automorphism together with ${\varepsilon}$ form a group isomorphic to ${\mathbb{Z}}_{2}$, which will be called the Weyl group of $m_{\alpha,\beta}$.
@@ 2.3 Casimir
Extend the isomorphisms ${\varepsilon}$ and $\pi_{n}$ defined in (2.21) and (2.22) to the algebra of operators on ${\cal Z}$. Then
${\varepsilon}({\mathcal{C}}_{\alpha,\beta})$ $={\mathcal{C}}_{\alpha,-\beta},$ (2.27)
@@ 2.4 Ladders
It is easy to see that if the representation of $m_{\alpha,\beta}$ on ${\cal Z}$ is irreducible and $N$ possesses an eigenvalue, then the spectrum of $N$ is a ladder. In fact this follows from
@@
For $n\in{\mathbb{C}}$ we define
${\cal Z}^{n}:=\{v\in{\cal Z}\ :\quad\quad Nv=nv\}.$ (2.30)
$A_{\pm}^{n\pm\frac{1}{2}}{\mathcal{C}}^{n}={\mathcal{C}}^{n\pm 1}A_{\pm}^{n\pm\frac{1}{2}}.$ (2.31)
@@ 2.5 Representation by 1st order differential operators
Thus it is a representation of Miller’s Lie algebra $m_{\alpha,\beta}$ with
@@ 2.6 Implementation of isomorphisms
Let us go back to ${\mathbb{C}}^{3}$ in the coordinates $s,t,z$. An implementation of the isomorphism $\pi_{n}$ defined in (2.22) is simple:
$\pi_{n}g(\cdot)=t^{-n}\cdot t^{n}.$ (2.42)
Let us now describe an implementation of the isomorphism ${\varepsilon}$. Let $\rho(z)$ solve
@@
$\begin{array}[]{l}Tg(w,z)=\rho(z)g(\frac{\sigma(z)}{w},z)\end{array}.$
@@
We easily see that in the coordinates (2.39), for all $F\in{\cal A}(\Omega)$,
$T^{n}F(z)=\rho(z)\sigma^{n}(z)F(z).$
@@ 3 Basic properties of hypergeometric class equations
In this section we introduce hypergeometric class equations. They will be presented as the eigenvalue equation of a certain second order differential operator ${\cal H}(\sigma,\kappa)$. We discuss a number of properties of these equations that can be described in a unified way: the basic symmetry, basic factorization and basic recurrence relations.
The operator ${\cal H}(\sigma,\kappa)$ is essentially the Casimir ${\mathcal{C}}$ of Miller’s Lie algebra at $N=0$, and all its properties discussed in this section follow from the properties of Miller’s Lie algebra analyzed in the previous section. However, this section can be read independently.
@@ 3.1 Remark about notation
Let $a,b,c$ be complex functions. Instead of saying that we consider the equation
@@ 3.2 Parametrization of hypergeometric class operators
where $\sigma,\tau,\eta$ satisfy the conditions (1.2). However, the parametrization of these equations with $\sigma,\tau,\eta$ is not always convenient. We will usually prefer to use a different parametrization, described below.
@@
${{\cal H}}(\sigma,\kappa)$ $:=\sigma(z)\partial_{z}^{2}+\big{(}\sigma^{\prime}(z)+\kappa(z)\big{)}\partial_{z}+\tfrac{1}{2}\kappa^{\prime}$ (3.5)
@@
${{\cal H}}(\sigma,\kappa)+\omega$ (3.6)
Note that ${{\cal H}}(\sigma,\kappa)$ coincides with ${{\mathcal{C}}}^{0}$, the Casimir operator for Miller’s Lie algebra restricted to the subspace $N=0$, expressed in the variable $z$, see (2.41).
@@ 3.3 Basic symmetry
${{\cal H}}(\sigma,\kappa)=\rho^{-1}(z)\partial_{z}\sigma(z)\rho(z)\partial_{z}+\tfrac{1}{2}\kappa^{\prime}.$ (3.8)
@@
$\begin{array}[]{l}\rho(z){{\cal H}}(\sigma,\kappa)\rho^{-1}(z)={{\cal H}}(\sigma,-\kappa).\end{array}$ (3.9)
@@
$\big{(}{{\cal H}}(\sigma,\kappa){+\omega\big{)}}F=0\ \Rightarrow\ \big{(}{{\cal H}}(\sigma,-\kappa){+\omega\big{)}}\rho F=0.$ (3.10)
Proof.  Using $\sigma(z)\partial_{z}\rho^{-1}(z)=-\kappa(z)\rho^{-1}(z)$, we obtain
$\rho(z){\cal H}(\sigma,\kappa)\rho^{-1}(z)$ $=\partial_{z}\sigma(z)\rho(z)\partial_{z}\rho^{-1}(z)+\frac{1}{2}\kappa^{\prime}$ (3.11)
$=\sigma(z)\partial_{z}^{2}+(\sigma^{\prime}(z)-\kappa(z))\partial_{z}-\frac{1}{2}\kappa^{\prime}$ $={\cal H}(\sigma,-\kappa).$ (3.12)
@@
Applying twice the basic symmetry we get the identity. Hence we obtain a group of symmetries of the hypergeometric class isomorphic to ${\mathbb{Z}}_{2}$.
Note that the basic symmetry of Theorem 3.10 corresponds to the symmetry ${\varepsilon}$ of Miller’s Lie algebra, see (2.21) and Theorem 2.1.